# STKDRec: Spatial-Temporal Knowledge Distillation for Takeaway Recommendation

### v1

@@ Abstract
The takeaway recommendation system is designed to recommend users’ future takeaway purchases based on their historical purchase behaviors, thereby improving user satisfaction and increasing merchant sales. Existing methods focus on incorporating auxiliary information or leveraging knowledge graphs to alleviate the sparsity issue of user purchase sequence data. However, two main challenges limit the performance of these approaches: (1) how to capture dynamic user preferences on complex geospatial information and (2) how to efficiently integrate spatial-temporal knowledge from graphs and sequence data with low calculation costs. In this paper, we propose a novel spatial-temporal knowledge distillation for takeaway recommendation model (STKDRec) based on the two-stage training process. Specifically, during the first pre-training stage, a spatial-temporal knowledge graph (STKG) encoder is pre-trained to extract the high-order spatial-temporal and collaborative associations within the STKG. During the second STKD stage, a spatial-temporal Transformer is employed to comprehensively model dynamic user preferences on various types of fine-grained geospatial information from a sequence perspective. Furthermore, the STKD strategy is introduced to adaptively fuse the rich spatial-temporal knowledge from the pre-trained STKG encoder and the spatial-temporal transformer while reducing the cost of model training. Extensive experiments on three real-world datasets show that our STKDRec significantly outperforms the state-of-the-art baselines. Our code is available at: https://github.com/Zhaoshuyuan0246/STKDRec.
@@ Introduction
Takeaway platforms, such as Yelp [^1] https://www.yelp.com/, Meituan [^2] https://www.meituan.com/, and Ele.me [^3] https://www.ele.me/, provide convenient online ordering and offline delivery services, playing an increasingly important role in people’s daily lives (Zhang et al. 2023). As the core service of these platforms, takeaway recommendation accurately recommends takeaways that align with user preferences based on their historical purchase behaviors. Such recommendation services enhance user satisfaction while increasing visibility and sales opportunities for merchants. In recent years, takeaway recommendations have gained significant attention from the research community and industry (Lin et al. 2022; Du et al. 2023a; Shi et al. 2024).
@@
Takeaway recommendation is essentially a sequential modeling task that aims to predict the user’s future takeaway purchases based on their historical purchase records. Existing methods model users’ purchase sequences by utilizing deep neural networks, such as recurrent neural networks (RNNs) and self-attention mechanisms. However, the sparsity issue arises since users often purchase only a few takeaways, limiting the performance of recommendations (Kang et al. 2023). Some studies attempt to incorporate various types of auxiliary information, such as brand, category, location, and area of interest (AOI), into purchase sequences. Some methods, such as KGDPL (Liu, Zhu, and Wu 2023), also utilize graph neural networks (GNNs) and Knowledge Graphs (KGs) (Wu et al. 2023; Chen et al. 2024) to explore higher-order user-takeaway relationships or rich semantics of takeaways, alleviating the data sparsity problem. Although these methods all achieve promising performance, two significant challenges still need to be addressed:
@@
(1) Failing to effectively capture dynamic user preferences on complex geospatial information. In takeaway recommendation scenarios, user preferences change dynamically over time and their current location. For example, users tend to purchase fast food at noon when at the workplace, while they prefer main meals in the evening when at home. However, complex geospatial information, including the distance between the user and the delivery merchant, the functional region of the merchant, and so on, as important factors affecting user preferences, has not been adequately explored. As illustrated in Figure 1, User #1 is primarily located in a workplace region during the day, frequently purchasing food that matches their preferences from nearby shops. However, when the user returns to the residential region in the evening, existing methods prioritize recommending nearby food candidates. Satisfying user preferences for foods from a more distant commercial region is neglected due to insufficient consideration of complex geospatial information. Therefore, modeling dynamic user preferences on complex geospatial information poses a challenge.
@@
(2) How to efficiently integrate spatial-temporal knowledge from graphs and sequence data with low calculation costs. The user’s purchase history is sequential data, while KGs are non-Euclidean structure data. Effectively integrating the advantages of these two types of heterogeneous data can alleviate the challenge of data sparsity and improve the accuracy of recommendations. However, due to the typically large scale of KGs, encoding them with GNNs significantly increases the parameter size and training overhead. Additionally, simple knowledge fusion methods, such as addition or concatenation, are not conducive to integrating these heterogeneous data for subsequent recommendations. Therefore, finding an effective method to fuse spatial-temporal knowledge from KGs and sequence data while reducing training overhead is crucial.
@@
To address these challenges, we propose a novel Spatial-Temporal Knowledge Distillation model for takeaway recommendation, termed STKDRec. STKDRec distills the offline teacher model’s knowledge of the graph structure to better enhance the student model’s ability to model users’ historical purchase sequences while reducing the training of parameters. STKDRec consists of two stages: pre-training and spatial-temporal knowledge distillation (STKD). During the pre-training stage, STKDRec introduces a spatial-temporal knowledge graph (STKG) and extracts rich spatial-temporal and collaborative associations between users and takeaways from STKG through pre-training an STKG encoder. During the STKD stage, a spatial-temporal transformer (ST-Transformer) is presented to capture dynamic user preferences on various types of fine-grained geospatial information from a sequential perspective. The spatial-temporal knowledge in the STKG encoder is transferred to the ST-Transformer through STKD, facilitating heterogeneous knowledge fusion and reducing the complexity of model training.
@@
  • We propose a novel spatial-temporal knowledge distillation model for takeaway recommendations, utilizing knowledge distillation to integrate spatial-temporal knowledge from STKG and sequence data, thereby addressing the sparsity issue in sequence data and effectively reducing training costs.
@@
Takeaway recommendation methods recommend personalized takeaways to users based on their historical purchase sequences. Early methods used Markov Chains (MC) (Rendle, Freudenthaler, and Schmidt-Thieme 2010) to model user sequences. With the development of deep learning, methods such as GNNs, RNNs, and self-attention mechanisms (Hidasi et al. 2015; Zhang et al. 2022; Chen et al. 2022; Shin et al. 2024) have been used to model users’ time-varying interests. Some methods attempt to capture users’ spatial-temporal interests by considering spatial-temporal information. StEN (Lin et al. 2022) models the spatial-temporal information of users and foods for click-through rate prediction in location-based service. BASM (Du et al. 2023a) integrates spatial-temporal features to capture user preferences at different times and locations. However, these methods primarily focus on simple spatial region information (e.g., geohash or AOI) without accounting for more fine-grained spatial information, such as spatial distance between the users and the takeaways. They fail to comprehensively and accurately capture dynamic user preferences in takeaway recommendations. Therefore, we propose a spatial-enhanced sequence representation encompassing various types of geospatial information and utilize an ST-Transformer to learn dynamic user preferences from a sequential perspective.
@@
Knowledge distillation (Hinton, Vinyals, and Dean 2015) achieves lightweight and performance enhancement by transferring knowledge from a teacher model to a student model. Existing methods (Kang et al. 2021; Xia et al. 2022; Zhu et al. 2021) can be categorized into distillation between models of the same type and distillation between models of different types. Distillation between models of the same type aims to reduce parameters, enabling the student model to approach the teacher model’s performance with fewer parameters. For example, TinyBERT (Devlin et al. 2018) uses the BERT model for pre-training and fine-tuning, distilling it into the smaller language model. Distillation can also be performed between different types of models. When there are structural or mechanistic differences between the models, the teacher model can provide a unique knowledge background and global understanding that are difficult for the student model to obtain independently. For instance, Graphless Neural Network (Zhang et al. 2021) translates the knowledge from the graph structure of GNNs into the MLP format. Motivated by these approaches, we leverage knowledge distillation to transfer the spatial-temporal knowledge in the STKG encoder to the ST-Transformer, thereby achieving the purpose of heterogeneous knowledge fusion and reducing the complexity of model training.
@@
The overall framework of STKDRec is shown in Figure 2, which consists of two stages: the pre-training stage and the STKD stage. During the pre-training stage, an STKG encoder is pre-trained to model the high-order spatial-temporal and collaboration associations between users and takeaways from the perspective of the graph structure. During the STKD stage, the ST-Transformer is employed to model the dynamic user preferences on complex spatial information from a sequential perspective. The STKD strategy facilitates the integration of spatial-temporal knowledge from STKG and sequence data.
@@
The STKG encoder, as the teacher model, is pre-trained to capture high-order spatial-temporal associations within the STKG, enabling it to model collaborative associations between users and takeaways from a graph structure perspective. Thus, we sample an STKG subgraph from STKG based on the user purchase sequence. An STKG encoder is used to aggregate spatial-temporal knowledge from the subgraph while integrating personalized user features.
@@
To model the higher-order spatial-temporal and collaborative associations between users and takeaways with $\mathcal{G}_{x}$, we utilize the GNNs to effectively encode the $\mathcal{G}_{x}$. In the $l$-th layer of the GNNs, the message passing and aggregation processes are defined as follows:
@@
To extract dynamic user preferences that vary across regions and distances from spatial-enhanced sequence representation, we present the spatial-temporal context attention mechanism. Specifically, this mechanism comprises $L$ layers of masked self-attention heads stacked together, transforming the input embedding $\hat{\text{E}}_{x}$ into the spatial-temporal context representation $\hat{\text{H}}_{x}\in\mathbb{R}^{n\times d}$. In each layer, we employ three independent linear transformation matrices $\text{W}_{i}^{q},\text{W}_{i}^{k},\text{W}_{i}^{v}\in\mathbb{R}^{d\times d^{\prime}}$ to transform the input embedding $\hat{\text{E}}_{x}$ into queries, keys, and values for the $i$-th scaled dot-product attention head, where $d^{\prime}=\frac{d}{K}$, and $i=1,2,\ldots,K$. The function is defined as follows:
@@
where $\hat{\text{h}}_{i}\in\mathbb{R}^{n\times d^{\prime}}$ denotes the output representation of the corresponding attention head. We concatenate the outputs from all attention heads to obtain the final spatial-temporal context representation $\hat{\text{H}}_{x}$.
$\hat{\text{H}}_{x}=\text{Concatenate}(\hat{\text{h}}_{1},\hat{\text{h}}_{2},\dots\hat{\text{h}}_{K})$ (8)
@@
Although the ST-Transformer captures dynamic user preferences from a sequential perspective but fails to capture the high-order spatial-temporal and collaborative associations between users and takeaways in the STKG. Thus, we propose the STKD strategy, which facilitates the transfer of spatial-temporal knowledge from the STKG encoder to the more efficient and lightweight ST-Transformer. This approach enables the heterogeneous fusion of spatial-temporal knowledge while significantly reducing the computational burden of training.
During the pre-training stage, the teacher model STKG encoder supervises the learning process using ground truth labels $\text{Y}_{x}\in\mathbb{R}^{1\times|\mathcal{V}|}$, generating soft labels $\text{Y}_{x}^{\prime}$ following pre-training. The pre-training loss is expressed as:
@@
During the STKD stage, our goal is to distill valuable spatial-temporal knowledge and collaborative associations from the STKG encoder, thereby enhancing the ST-Transformer’s ability to capture user preferences from different perspectives and promoting more efficient, streamlined learning. To achieve this, we train the student model, ST-Transformer, to emulate the soft labels of the teacher model, effectively transferring spatial-temporal knowledge from the teacher to the student. The distillation loss is defined as:
@@
We select three publicly available city takeaway recommendation datasets for evaluation: Wuhan, Sanya, and Taiyuan, which are provided by the well-known takeaway platform Ele.me [^4] The dataset is available at https://tianchi.aliyun.com/dataset/131047.. Each dataset contains users’ purchase sequences and associated attributes. The attributes mainly include geospatial features, such as the geohash6 and AOI, where users and takeaways are located, temporal features, such as the timestamps and weekdays when users purchase takeaway, and additional user and takeaway attributes. To ensure the quality of data, we apply the following processing steps. Initially, we remove the users and takeaways with empty geohash6 attribute values to clean the dataset. Secondly, we retain features from the cleaned dataset and calculate the spherical distance as a spatial distance feature based on the geohash6 values of users and takeaways. Finally, for each user purchase sequence, the last purchased takeaway is designated as test data, the second-to-last as validation data, and the rest as training data. A statistical summary of the processed datasets is shown in Table 1.
@@
To evaluate the effectiveness of our model, we compare it with the following eight representative baselines: Caser (Tang and Wang 2018), GRU4Rec (Hidasi et al. 2015), SASRec (Kang and McAuley 2018), BERT4Rec (Sun et al. 2019), DuoRec (Qiu et al. 2022), FEARec (Du et al. 2023b), GCL4SR (Zhang et al. 2022), MAERec (Ye, Xia, and Huang 2023), and BSARec (Shin et al. 2024).
@@
All evaluation methods are implemented in PyTorch. The hyper-parameters for these methods are chosen according to the original papers, and the optimal settings are selected based on model performance on the validation data. For STKDRec, we conduct experiments with the following hyper-parameters: For the STKG encoder, the sampling depth $m$ is set to 2, and the number of neighbor nodes sampled $s$ is selected from {[5,5], [10,10], [15,15], [20,20]}. For the ST-Transformer module, we set the number of self-attention blocks and attention heads to 2 and the embedding dimension to 256. For STKD, the temperature $\tau$ is selected from {1, 3, 5, 7, 9}, and the $\alpha$ is fixed at 0.2. We use Adam as the optimizer, with the learning rate, $\beta_{1}$, and $\beta_{2}$ set to 0.001, 0.9, and 0.98, respectively. The batch size and maximum sequence length $n$ are both set to 128. An early stopping strategy is applied based on the performance of the validation data. The implementation is carried out in PyTorch on a single NVIDIA A40 GPU with 48 GB of memory.
@@
Table 2 presents the overall experimental results of STKDRec and baselines. From Table 2, we have the following vital observations: STKDRec achieves the best performance across all datasets compared to all baselines, highlighting the effectiveness of our STKDRec. The second-best model BSARec considers fine-grained user sequential patterns without accounting for spatial-temporal and collaborative associations between users and takeaways. These limitations result in BSARec performing worse than STKDRec.
Graph-based methods GCL4Rec and MAERec focus on user-item graphs to capture the collaborative association between users and takeaways but lack the integration of temporal, spatial, and additional auxiliary information, limiting their ability to capture users’ preferences over time and geospatial information. In takeaway recommendation scenarios, these factors significantly influence users’ purchase behavior. This is why GCL4SR and MAERec perform worse than STKDRec. Feature-enhanced methods DuoRec and FeaRec rely on static features or limited external knowledge, making it difficult to effectively capture the dynamic changes in user preferences. Our STKDRec captures dynamic user preferences on complex geospatial information and integrates the spatial-temporal and collaborative associations within the STKG.
@@
Figure 3: Study on different knowledge fusion methods. multi refers to multiplication, cat refers to concatenation, add refers to addition, STKD denotes our proposed model, and Time indicates the model training duration.
@@
To assess the contribution of each component of STKDRec, we perform ablation studies on all datasets. The results of the STKDRec variants are shown in Table 3. -w/o SP denotes a variant without spatial position embedding. -w/o F and -w/o C denotes two variants considering only spatial region and spatial distance information separately. -w/o KD denotes a variant using only the ST-Transformer module; -w/o SP + KD denotes a variant using only the original Transformer to encode sequence. The results in Table 3 show that the variants -w/o SP, -w/o F, and -w/o C consistently perform worse than STKDRec on all datasets, highlighting the importance of geospatial information for modeling dynamic user preferences in takeaway recommendations. The results for -w/o KD further indicate that the STKD strategy is crucial for facilitating the fusion of spatial-temporal knowledge from STKG and sequence data. The significant drop in performance for -w/o KD+SP demonstrates the effectiveness of all the proposed improvements.
To validate the efficacy of the STKD strategy, we replace it with other knowledge fusion strategies, including concatenation, addition, and multiplication. The results are shown in Figure 3. STKDRec achieves superior performance and better training efficiency than other variants. This is because the knowledge distributions of the STKG encoder and the ST-Transformer differ significantly and exist in distinct vector spaces. Simple combing them fails to effectively integrate the differing distributions and may transform the distribution of the STKG encoder into noise, thereby damaging the distribution of the ST-Transformer. In addition, these variants necessitate additional computations to align the distinct vector spaces, which further compromises the overall efficiency of the model. STKD directly aligns the predicted distributions of the ST-Transformer and STKG encoder, facilitating smoother and more effective knowledge transfer.
@@
We provide a case study to explore the prediction results of STKDRec and the variant -w/o KD. Given the historical purchase sequence of user u420, Figure 5 illustrates the takeaway recommendation rankings and corresponding probabilities for u420 at 4:40 PM on Wednesday at their workplace. Coffee and donuts have strong spatial-temporal and collaborative associations, as they are frequently purchased together by users in workplace areas during afternoon tea time. The variant -w/o KD only considers the foods previously purchased by the user, failing to capture these collaborative associations, which leads to incorrect recommendations. However, by incorporating STKG, STKDRec can capture the spatial-temporal and collaborative relationships between coffee and donuts, allowing it to accurately recommend the donut for u420. This case further demonstrates the effectiveness of integrating the spatial-temporal knowledge and collaborative associations between the STKG and sequence data.

### v2

# Spatial-Temporal Knowledge Distillation for Takeaway Recommendation

@@ Abstract
The takeaway recommendation system aims to recommend users’ future takeaway purchases based on their historical purchase behaviors, thereby improving user satisfaction and boosting merchant sales. Existing methods focus on incorporating auxiliary information or leveraging knowledge graphs to alleviate the sparsity issue of user purchase sequences. However, two main challenges limit the performance of these approaches: (1) capturing dynamic user preferences on complex geospatial information and (2) efficiently integrating spatial-temporal knowledge from both graphs and sequence data with low computational costs. In this paper, we propose a novel spatial-temporal knowledge distillation model for takeaway recommendation (STKDRec) based on the two-stage training process. Specifically, during the first pre-training stage, a spatial-temporal knowledge graph (STKG) encoder is trained to extract high-order spatial-temporal dependencies and collaborative associations from the STKG. During the second spatial-temporal knowledge distillation (STKD) stage, a spatial-temporal Transformer (ST-Transformer) is employed to comprehensively model dynamic user preferences on various types of fine-grained geospatial information from a sequential perspective. Furthermore, the STKD strategy is introduced to transfer graph-based spatial-temporal knowledge to the ST-Transformer, facilitating the adaptive fusion of rich knowledge derived from both the STKG and sequence data while reducing computational overhead. Extensive experiments on three real-world datasets show that STKDRec significantly outperforms the state-of-the-art baselines.
Codehttps://github.com/Zhaoshuyuan0246/STKDRec
@@ Introduction
Takeaway platforms, such as Yelp, Meituan, and Ele.me, provide convenient online ordering and offline delivery services, playing an increasingly important role in people’s daily lives (Zhang et al. 2023). As the core service of these platforms, takeaway recommendation aim to accurately recommend takeaways that align with user preferences based on their historical purchase behaviors. Such recommendation services enhance user satisfaction while increasing visibility and sales opportunities for merchants. In recent years, takeaway recommendations have gained significant attention from the research community and industry (Lin et al. 2022; Du et al. 2023a; Shi et al. 2024).
@@
Takeaway recommendation is essentially a sequential modeling task that aims to predict the user’s future takeaway purchases based on their historical purchase records. Existing methods model these purchase sequences by utilizing deep neural networks, such as recurrent neural networks (RNNs) and self-attention mechanisms. However, the sparsity issue arises since users often purchase only a few takeaways, limiting the performance of recommendations (Kang et al. 2023). Some studies attempt to incorporate various types of auxiliary information, such as brand, category, location, and area of interest (AOI), into purchase sequences. Some methods, such as KGDPL (Liu, Zhu, and Wu 2023), also utilize graph neural networks (GNNs) and Knowledge Graphs (KGs) (Wu et al. 2023; Chen et al. 2024) to explore higher-order user-takeaway relationships or rich semantics of takeaways, alleviating the data sparsity problem. Although these methods all achieve promising performance, two significant challenges still need to be addressed:
@@
(1) Failing to effectively capture dynamic user preferences on complex geospatial information. In takeaway recommendation scenarios, user preferences change dynamically over time and their current location. For example, users tend to purchase fast food at noon when at the workplace, while they prefer main meals in the evening when at home. However, complex geospatial information, including the distance between the user and the delivery merchant, the functional region of the merchant, and so on, as important factors affecting user preferences, has not been adequately explored. As illustrated in Figure 1, User #1 is primarily located in a workplace region during the day, frequently purchasing food that matches their preferences from nearby shops. However, when the user returns to the residential region in the evening, existing methods prioritize recommending nearby food candidates. Satisfying user preferences for foods from a more distant commercial region is neglected due to insufficient consideration of complex geospatial information. Therefore, capturing dynamic user preferences on complex geospatial information poses a challenge.
@@
(2) How to efficiently integrate spatial-temporal knowledge from both graphs and sequence data with low computational costs. The user’s purchase history is sequential data, while KGs are non-Euclidean structure data. Effectively integrating the advantages of these two types of heterogeneous data can alleviate the challenge of data sparsity and improve the accuracy of recommendations. However, due to the typically large scale of KGs, encoding them with GNNs significantly increases computational overhead. Additionally, simple knowledge fusion methods, such as addition or concatenation, are not conducive to integrating these heterogeneous data for subsequent recommendations. Therefore, finding an effective method to fuse spatial-temporal knowledge from both graphs and sequence data while reducing computational costs is crucial.
@@
To address these challenges, we propose a novel Spatial-Temporal Knowledge Distillation model for takeaway recommendation, termed STKDRec. The model distills the offline teacher model’s knowledge of the graph structure to better enhance the student model’s ability to model users’ historical purchase sequences while improving computational efficiency. STKDRec consists of two stages: pre-training and spatial-temporal knowledge distillation (STKD). During the pre-training stage, STKDRec constructs a spatial-temporal knowledge graph (STKG) and extracts high-order spatial-temporal dependencies and collaborative associations between users and takeaways from STKG through training an STKG encoder. During the STKD stage, a spatial-temporal Transformer (ST-Transformer) is presented to capture dynamic user preferences on various types of fine-grained geospatial information from a sequential perspective. Through the STKD strategy, graph-based spatial-temporal knowledge from the STKG encoder is effectively transferred to the ST-Transformer, facilitating heterogeneous knowledge fusion with low computational costs.
@@
  • We propose a novel spatial-temporal knowledge distillation model for takeaway recommendations, utilizing knowledge distillation to fuse spatial-temporal knowledge from both STKG and sequence data, thereby addressing the sparsity issue in sequence data and effectively reducing computational overhead.
@@
Takeaway recommendation methods recommend personalized takeaways to users based on their historical purchase sequences. Early methods used Markov Chains (MC) (Rendle, Freudenthaler, and Schmidt-Thieme 2010) to model user sequences. With the development of deep learning, methods such as GNNs, RNNs, and self-attention mechanisms (Hidasi et al. 2015; Zhang et al. 2022; Chen et al. 2022; Shin et al. 2024) have been used to model users’ time-varying interests. Some methods attempt to capture users’ spatial-temporal interests by considering spatial-temporal information. StEN (Lin et al. 2022) models the spatial-temporal information of users and foods for click-through rate prediction in location-based service. BASM (Du et al. 2023a) integrates spatial-temporal features to capture user preferences at different times and locations. However, these methods primarily focus on spatial region information (e.g., geohash or AOI) without accounting for fine-grained spatial information of different types, such as spatial distance between the users and the takeaways. They fail to comprehensively and accurately capture dynamic user preferences in takeaway recommendations. Therefore, we propose a spatial-enhanced sequence representation encompassing various types of geospatial information and utilize an ST-Transformer to learn dynamic user preferences from a sequential perspective.
@@
Knowledge distillation (Hinton, Vinyals, and Dean 2015) achieves lightweight and performance enhancement by transferring knowledge from a teacher model to a student model. Existing methods (Kang et al. 2021; Xia et al. 2022; Zhu et al. 2021) can be categorized into distillation between models of the same type and distillation between models of different types. Distillation between models of the same type aims to reduce parameters, enabling the student model to approach the teacher model’s performance with fewer parameters. For example, TinyBERT (Devlin et al. 2018) uses the BERT model for pre-training and fine-tuning, distilling it into the smaller language model. Distillation can also be performed between different types of models. When there are structural or mechanistic differences between the models, the teacher model can provide a unique knowledge background and global understanding that are difficult for the student model to obtain independently. For instance, Graphless Neural Network (Zhang et al. 2021) translates the knowledge from the graph structure of GNNs into the MLP format. Motivated by these approaches, we leverage knowledge distillation to transfer the graph-based spatial-temporal knowledge from the STKG encoder to the ST-Transformer, thereby achieving the purpose of heterogeneous knowledge fusion and reducing computational costs.
@@
The overall framework of STKDRec is shown in Figure 2, which consists of two stages: the pre-training stage and the STKD stage. During the pre-training stage, an STKG encoder is trained to model the high-order spatial-temporal dependencies and collaborative associations between users and takeaways from the perspective of the graph structure. During the STKD stage, an ST-Transformer is employed to model the dynamic user preferences on complex spatial information from a sequential perspective. The STKD strategy facilitates the fusion of spatial-temporal knowledge from both STKG and sequence data.
@@
The STKG encoder, as the teacher model, is pre-trained to capture high-order spatial-temporal knowledge from the STKG. Thus, we sample an STKG subgraph from STKG based on the user purchase sequence. The STKG encoder is used to aggregate spatial-temporal knowledge from the subgraph while integrating personalized user features.
@@
To model the higher-order spatial-temporal dependencies and collaborative associations between users and takeaways with $\mathcal{G}_{x}$, we utilize the GNNs to effectively encode the $\mathcal{G}_{x}$. In the $l$-th layer of the GNNs, the message passing and aggregation processes are defined as follows:
@@
To extract dynamic user preferences that vary across regions and distances from spatial-enhanced sequence representation, we present the spatial-temporal context attention mechanism. Specifically, this mechanism comprises $L$ layers of mask self-attention layers stacked together, transforming the input embedding $\hat{\text{E}}_{x}$ into the spatial-temporal context representation $\hat{\text{H}}_{x}\in\mathbb{R}^{n\times d}$. In each layer, we employ three independent linear transformation matrices $\text{W}_{i}^{q},\text{W}_{i}^{k},\text{W}_{i}^{v}\in\mathbb{R}^{d\times d^{\prime}}$ to transform the input embedding $\hat{\text{E}}_{x}$ into queries, keys, and values for the $i$-th scaled dot-product attention head, where $d^{\prime}=\frac{d}{K}$, and $i=1,2,\ldots,K$. The function is defined as follows:
@@
where $\hat{\text{h}}_{i}\in\mathbb{R}^{n\times d^{\prime}}$ denotes the output representation of the corresponding attention head. We concatenate the outputs from all attention heads to obtain the final spatial-temporal context representation $\hat{\text{H}}_{x}$,
$\hat{\text{H}}_{x}=\text{FFN}(\text{Concatenate}(\hat{\text{h}}_{1},\hat{\text{h}}_{2},\dots\hat{\text{h}}_{K})),$ (8)
where FFN(·) denotes the feed-forward network.
@@
Although the ST-Transformer captures dynamic user preferences from a sequential perspective, it fails to capture the high-order spatial-temporal dependencies and collaborative associations between users and takeaways from the STKG. Thus, we propose the STKD strategy, which facilitates the transfer of graph-based spatial-temporal knowledge from the STKG encoder to the more efficient and lightweight ST-Transformer. This approach enables the heterogeneous fusion of spatial-temporal knowledge while significantly reducing the computational costs.
During the pre-training stage, the teacher model STKG encoder supervises the learning process using ground truth labels $\text{Y}_{x}\in\mathbb{R}^{1\times|\mathcal{V}|}$, generating soft labels $\text{Y}_{x}^{\prime}$ following pre-training. The pre-training loss is defined as:
@@
During the STKD stage, our goal is to distill valuable spatial-temporal knowledge from the STKG encoder, thereby enhancing the ST-Transformer’s ability to capture user preferences from different perspectives and promoting more efficient, streamlined learning. To achieve this, we train the student model, ST-Transformer, to emulate the soft labels of the teacher model, effectively transferring knowledge from the teacher to the student. The distillation loss is defined as:
@@
We select three publicly available city takeaway recommendation datasets for evaluation: Wuhan, Sanya, and Taiyuan, which are provided by the well-known takeaway platform Ele.me. Each dataset contains users’ purchase sequences and associated attributes. The attributes mainly include geospatial features, such as the geohash6 and AOI, where users and takeaways are located, temporal features, such as the timestamps and weekdays when users purchase takeaways, and additional user and takeaway attributes. To ensure the quality of data, we apply the following processing steps. Initially, we remove the users and takeaways with empty geohash6 attribute values to clean the dataset. Secondly, we retain features from the cleaned dataset and calculate the spherical distance as a spatial distance feature based on the geohash6 values of users and takeaways. Finally, for each user purchase sequence, the last purchased takeaway is designated as test data, the second-to-last as validation data, and the rest as training data. A statistical summary of the processed datasets is shown in Table 1.
@@
To evaluate the effectiveness of our model, we compare it with the following nine representative baselines: Caser (Tang and Wang 2018), GRU4Rec (Hidasi et al. 2015), SASRec (Kang and McAuley 2018), BERT4Rec (Sun et al. 2019), DuoRec (Qiu et al. 2022), FEARec (Du et al. 2023b), GCL4SR (Zhang et al. 2022), MAERec (Ye, Xia, and Huang 2023), and BSARec (Shin et al. 2024).
@@
All evaluation methods are implemented in PyTorch. The hyper-parameters for these methods are chosen according to the original papers, and the optimal settings are selected based on model performance on the validation data. For STKDRec, we conduct experiments with the following hyper-parameters. For the STKG encoder, the sampling depth $m$ is set to 2, and the number of neighbor nodes sampled $s$ is selected from {[5,5], [10,10], [15,15], [20,20]}. For the ST-Transformer module, we set the number of self-attention blocks and attention heads to 2 and the embedding dimension to 256. For STKD, the temperature $\tau$ is selected from {1, 3, 5, 7, 9}, and the $\alpha$ is fixed at 0.2. We use Adam as the optimizer, with the learning rate, $\beta_{1}$, and $\beta_{2}$ set to 0.001, 0.9, and 0.98, respectively. The batch size and maximum sequence length $n$ are both set to 128. An early stopping strategy is applied based on the performance of the validation data. The implementation is carried out in PyTorch on a single NVIDIA A40 GPU with 48GB of memory.
@@
Table 2 presents the overall experimental results of STKDRec and baselines. We have the following vital observations: STKDRec achieves the best performance across all datasets compared to all baselines, highlighting the effectiveness of our STKDRec. The second-best model BSARec considers fine-grained user sequential patterns without accounting for spatial-temporal dependencies and collaborative associations between users and takeaways. These limitations result in BSARec performing worse than STKDRec.
Graph-based methods GCL4Rec and MAERec focus on user-item graphs to capture the collaborative association between users and takeaways but lack the integration of temporal, spatial, and additional auxiliary information, limiting their ability to capture users’ preferences over time and geospatial information. In takeaway recommendation scenarios, these factors significantly influence users’ purchase behavior. This is why GCL4SR and MAERec perform worse than STKDRec. Feature-enhanced methods DuoRec and FeaRec rely on static features or limited external knowledge, making it difficult to effectively capture the dynamic changes in user preferences. Our STKDRec captures dynamic user preferences on complex geospatial information and integrates the spatial-temporal knowledge within the STKG.
@@
Figure 3: Study on different knowledge fusion methods. Multi refers to multiplication, Cat refers to concatenation, Add refers to addition, STKD denotes our proposed strategy, and Time indicates model training and prediction duration.
@@
To assess the contribution of each component of STKDRec, we perform ablation studies on all datasets. The results of the STKDRec variants are shown in Table 3. -w/o SP denotes a variant without spatial position embedding. -w/o F and -w/o C denote two variants considering only spatial region and spatial distance information separately. -w/o KD denotes a variant using only the ST-Transformer module; -w/o SP + KD denotes a variant using only the original Transformer to encode sequence. The results in Table 3 show that the variants -w/o SP, -w/o F, and -w/o C consistently perform worse than STKDRec on all datasets, highlighting the importance of geospatial information for modeling dynamic user preferences in takeaway recommendations. The results for -w/o KD further indicate that the STKD strategy is crucial for facilitating the fusion of spatial-temporal knowledge from STKG and sequence data. The significant drop in performance for -w/o KD+SP demonstrates the effectiveness of all the proposed improvements.
To validate the efficacy of the STKD strategy, we replace it with other knowledge fusion strategies, including concatenation, addition, and multiplication. The results are shown in Figure 3. STKDRec achieves superior performance and better training and prediction efficiency than other variants. This is because the knowledge distributions of the STKG encoder and the ST-Transformer differ significantly and exist in distinct vector spaces. Simple combining them fails to effectively integrate the differing distributions and may transform the distribution of the STKG encoder into noise, thereby damaging the distribution of the ST-Transformer. In addition, these variants necessitate additional computations to align the distinct vector spaces, which further compromises the overall efficiency of the model. STKD directly aligns the predicted distributions of the ST-Transformer and STKG encoder, facilitating smoother and more effective knowledge transfer.
@@
To evaluate the impact of the STKD strategy, we conduct a case study comparing the recommendation results of STKDRec and its variant -w/o KD. Given user u420’s historical purchase sequence, Figure 5 illustrates the Top-5 takeaway recommendation rankings and their corresponding probabilities for u420 at 4:40 PM on Wednesday at their workplace. Coffee and donuts exhibit strong spatial-temporal dependencies and frequent co-purchasing behaviors, driven by their recurrent joint consumption in workplace areas during afternoon tea time. The variant -w/o KD only relies on the user’s historical purchasing data, failing to account for these collaborative associations, which leads to incorrect recommendations. However, by incorporating STKG, STKDRec can capture the spatial-temporal and collaborative relationships between coffee and donuts, allowing it to recommend the donuts for u420 accurately. This case further demonstrates the effectiveness of utilizing the STKD strategy to integrate the spatial-temporal knowledge from both STKG and sequence data.